5z^2-33z-14 Factor.
Answer by edjones(8007)

5z^2-33z-14
5*14=-70 The 2 factors of -70 whose sum is -33 are -35 and +2.
5z^2+2z-35z-14
=z(5z+2)-7(5z+2) factoring by grouping
=(z-7)(5z+2)
.
Ed